Q1.> An electron is movin in XY- plane , makin an angle relative to the position
X-axis.The specific charge of electron is 1.758 x 10 par11 C kg-1 . A uniform magnetic field of induction 0.8 Wbm-2 acts along the positive X-axis . Then the particle will pass periodically through the X-axis with a period
 
 
a> 1.406x10par 11 s              b> 2.198x10par -11 s
 
 
c> 4.465x10par-11 s              d>0.711x10par-11 s
 
 
tell me the ans


Grade:9

1 Answers

ROSHAN MUJEEB
3 years ago
Magnetic force is given asF=qvBsinθit is given that angle between fieldBand velocityvis900andq=1.6×10−19Cso we get the value of force asF=1.6×10−16Nsom=1is the required answer.